OpenStudy (anonymous):

^-^ Help is very much needed, math isn't my strong suit. What is the equation of the line expressed in slope-intercept form? A. y = 4x – 2 B. y = –2x + 4 C. y = 2x + 2 D. y = –2x – 4

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

slop = y=mx+b wheres the line?

pooja195 (pooja195):

ok where is the line hitting on the y axis?

OpenStudy (anonymous):

Uh, -4 I think

pooja195 (pooja195):

yes so we have our y intercept now what about x where is the line hitting?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

good so -4 is the y intercept and -2 is the x intercept now plug it in! y=mx+b

OpenStudy (anonymous):

I'm confused :/

pooja195 (pooja195):

ok X intercept is -2 where does it fit in y=mx+b?

OpenStudy (anonymous):

Uhhh... y=m-2+b??

pooja195 (pooja195):

nope your making it harder than it needs to be ok\[y=mx+b\] this is the slop intercept equation got it?

OpenStudy (anonymous):

Okay.

pooja195 (pooja195):

so X intercept which is -2 you would plug in place of M so what does your equation look like? once you plug in the numbers?

OpenStudy (anonymous):

y=-2x+4

pooja195 (pooja195):

omg your so close!!! remeber its -4 try again :) you're almost there

pooja195 (pooja195):

do you know?

OpenStudy (anonymous):

y=-2x+-4

OpenStudy (anonymous):

Is it D?

pooja195 (pooja195):

take out the + sign and yep! well done!! :D

OpenStudy (anonymous):

Thank you again!

pooja195 (pooja195):

Hahah no probs :)

OpenStudy (ashleyisakitty):

For the reference: Slope-intercept:\[y=mx+b\] Point slope: \[y-y _{1}=m(x-x _{1})\] RAO or slope: \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]